Please help me solve this equation (step by step)
Using Elimination
x+y=1
-2x+y=4

Answer by mananth(16946)

1 x + 1 y = 1 .............1
-2 x + 1 y = 4 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -1
1 x 1 y = 1
2 x -1 y = -4
Add the two equations
3 x = -3
/ 3
x = -1
plug value of x in (1)
1 x + 1 y = 1
-1 + 1 y = 1
1 y = 1+1
1 y = 2
y = 2